the areas of three ajacent faces of cuboids are 108cm,72cm and 96 respectively. find the volume of the cuboid ,also find the dimensions of the cuboid
 
 

Grade:9

1 Answers

Naman Agarwal
7 years ago
Let length breadth Nd hieght of a cuboid be x,y,z.. Multiply the area of three adjacent faces of cuboid Xy × yz × zx= 108×72×96(Xyz)^2= 746496Xyz= √746496= 864Xyz/yz= 864/72X= 12Xy= 96. So y= 8Xz= 108. So z=9Ans....So dimensions are 12,9,8Nd volume is 864....